# A transiting multi-planet system in the 61 million year old association Theia 116

### v1

@@ 2.3 Ground-based Photometry
We obtained ground-based photometry as part of the TESS Follow-up Observing Program (TFOP; Collins, 2019) [^1] https://tess.mit.edu/followup. We used the TESS Transit Finder, which is a customized version of the Tapir software package (Jensen, 2013), to schedule our transit observations. A summary of our ground-based photometric observations is presented in Table 1, with the light curves shown in Figure 4. The ground-based observations confirm the transits of TIC 434398831 b and c are on target to within 3.1”, and clear the field for nearby eclipsing binary (NEB) scenarios $<2.5$’. We made use of the raw ground-based photometry and simultaneously detrended the transit observations in our global model.
@@ 2.4 Reconnasance Spectroscopy
We obtained a set of spectroscopic observations to characterize the spectroscopic properties of the host star and to search for signatures of astrophysical false positive scenarios. We obtained seven epochs of observations with the Tillinghast Reflector Echelle Spectrograph (TRES; Fűrész 2008) on the 1.5 m reflector at the Fred Lawrence Whipple Observatory, Arizona. TRES is a high-resolution echelle spectrograph with a resolving power of $R\approx 44,000$, with a wavelength coverage of $3850-9100\,$Å. Observing strategy and spectral extraction techniques are outlined in Buchhave et al. (2012), spectroscopic atmospheric parameters are derived from each spectrum as per Buchhave et al. (2010), and velocities were derived from a cross-correlation analysis over multiple spectral orders as per Quinn et al. (2012), and are presented in Table 2.
@@ 3.1.1 Kinematics
Figure 5: Color-Magnitude Diagram (CMD) for candidate Theia 116 members and TIC 434398831 (red star). We plot candidate members with measured rotation periods in orange, filled circles. Candidates without measured rotations are plotted as empty circles. We plot the Kepler Input Catalog (KIC; Brown et al., 2011) CMD for reference (grey).
@@
TIC 434398831 was identified as a candidate member of Theia 116 in Kounkel & Covey (2019). In brief, Kounkel & Covey (2019) performed a search for comoving stellar populations through analysis of spatial motions from Gaia DR2 (Gaia Collaboration et al., 2018b) via the unsupervised learning algorithm HDBSCAN (McInnes et al., 2017). Kounkel & Covey (2019) identified 1108 candidate members of Theia 116. We used Gaia DR3 (Gaia Collaboration et al., 2021) to determine Theia 116 candidate members’ spatial positions and velocities. Figure 6 shows TIC 434398831 relative to the space motion distribution of the Theia 116 group.
@@ 3.1.2 Stellar Rotation
Figure 7: Top panel: Lomb-Scargle periodogram for TIC 434398831 both the full TESS light curve (navy) and each sector of TESS observations (grey). We measure a stellar rotation period of $P_{\mathrm{rot}}=1.75\pm 0.12$ days.
@@
Figure 8: Upper panel: Rotation periods for rotationally validated Theia 116 members (orange circles) and TIC 434398831 (red star) as a function of stellar effective temperature. We plot measured rotations of well-characterized clusters for comparison– 40 Myr Tuc-Hor (left, Popinchalk et al., 2023), 80 Myr $\alpha$ Per (middle, Boyle & Bouma, 2023), and 120 Myr Pleiades (right, Rebull et al., 2016). Lower panel: Lithium 6708Å  doublet equivalent widths (EW) of a subsample of Theia 116 members (orange circles) and TIC 434398831 (red circle). Rotationally validated members of Theia 116 are the filled circles, while empty circles do not exhibit photometric signatures of youth, and are likely contaminating field stars. Literature Li EW’s of Tuc-Hor (Popinchalk et al., 2023), $\alpha$ Per (Balachandran et al., 2011), and Pleiades (Bouvier et al., 2018) members are plotted for comparison. The spread in both rotation and Li abundances of Theia 116 members are consistent with youth. The rotation spread is more consistent with spreads of the pre-zero age main sequence (ZAMS) clusters Tuc-Hor and $\alpha$ Per, than the spread of the ZAMS Pleiades cluster.
@@
We visually examined the full light curve of each candidate member for (1) rotational variability, (2) binarity, (3) non-rotation variability, or (4) no variability. We checked for aliasing of the true rotation period by phase-folding the light curve at the measured Lomb-Scargle period, half, and double the period. We were able to rotationally validate 224 members of Theia 116, including TIC 434398831, for which we measured a rotation period of $P_{\mathrm{rot}}=1.75\pm 0.12$ days (see Figure 7). Our measured rotation periods for Theia 116 members are shown in Figure 8.
@@ 3.1.3 Lithium Depletion
We measured the equivalent width of the lithium doublet at 6708 Å as per Zhou et al. (2021). We first performed a least-squares deconvolution analysis to derive the line profile of each star, and thereby measure its rotational velocity. We then fit the spectral region around the 6708 Å line via a set of Gaussian profiles. We assumed the two lines of the lithium doublet have the same amplitude and the same width as that of the stellar rotational broadening velocity. We also simultaneously modeled the nearby Fe I line at 6707.43 Å to account for its contamination to the lithium doublet line strength, adopting the same broadening width, but allowing its amplitude to be free. We adopted the integral below our model of the lithium doublet as the equivalent width of the lithium line. All spectra were checked for signatures of binarity, no members exhibited double-line profiles. The measured lithium abundances are presented in Table 4. We compare our measured lithium equivalent widths against literature values for well-characterized clusters in Figure 8.
@@ 3.1.4 Isochrone modeling of membership SEDs
We fitted for the cluster age, extinction, and stellar mass of each rotationally validated member. These parameters were then interpolated onto a MIST isochrone (Dotter, 2016) with minimint (Koposov, 2021) to derive estimated stellar magnitudes. We assumed solar metallicity and Gaia parallax values. We compared estimated magnitudes to observed magnitudes from the Gaia DR3 G, BP, and RP bands, WISE W1, W2, and W3 bands, and 2MASS J, H, and Ks bands. The SED of TIC 434398831 is presented in Figure 9. The best-fit cluster age and its associated uncertainties were explored via an MCMC implemented in emcee. We estimated a SED cluster age of 55$\pm$4 Myr.
@@ 3.1.5 Age of Theia 116
We attempted numerous joint age analyses that incorporated lithium distributions, isochrones, and rotation distributions together for Theia 116, but derived age posteriors with uncertainties too small to be realistic (69±$\pm 3$ Myr), likely resulting from intrinsic uncertainties in stellar models unaccounted for in our analysis. For the remainder of this work, we therefore adopt a more conservative cluster age of $\pm 6$ Myr, which encompasses the posteriors of the age analyses from independent lithium distribution modeling and the isochrone modeling.
@@ 3.2 Stellar flare rate of TIC 434398831
In addition to heightened photometric variability induced by rapid rotation, stellar flares from active young stars bring additional challenges when searching for transit planets (e.g. AU Mic c Plavchan et al., 2020). At $\mathord{\sim}61$ Myr, young G-type stars are still expected to have a heightened stellar flare rate when compared to their gigayear old counterparts (Feinstein et al., 2020). Visual examination of the TESS light curves showed evidence for stellar flares detected in the 10 min FFI observations. However, most flare identification algorithms are tuned for TESS 20 second or 2 minute cadence data (e.g. Feinstein et al., 2020).
[…]
@@ 5 Investigating possible false-positive scenarios
Due to the large size of TESS’s pixels, some transit signals that pass vetting procedures may result from false positives. As TIC 434398831 b and c were not released as a TOI, we investigate possible instrumental and astrophysical false-positive scenarios to validate the planets. The transit detections of both TIC 434398831 b and c by multiple instruments ruled out instrumental artifact false-positive scenarios. Spectroscopic follow-up observations ruled out eclipsing binary scenarios.
@@
where $t_{12}$ is the ingress duration, $t_{13}$ is the time between first and third contact, and $\delta$ is the transit depth. An unresolved star would have to be within $\Delta m_{\mathrm{\textit{TESS}}}\lesssim 3.16$ and $\Delta m_{\mathrm{\textit{TESS}}}\lesssim 2.48$ (at a $3\sigma$ upper limit) to produce the signals corresponding to the transits of TIC 434398831 b and c respectively. We estimated the density of stars within $\Delta m_{\mathrm{\textit{TESS}}}\leq 3.16$ in the general area of TIC 434398831 to determine the likelihood of an unresolved background star being the source of the transit signals. We derive a density of $8.84\times 10^{-5}$ stars per square arcsecond within $\Delta m_{\mathrm{\textit{TESS}}}\leq 3.16$ of TIC 434398831.
We also made use of our TRES spectroscopic observations to help establish the magnitude limits of hypothetical stellar neighbors that may be blended with our target star. Following Zhou et al. (2021), we used the line profiles derived from our TRES observations to determine if certain stellar companion scenarios can be ruled out, as they would yield a detectable set of lines in the spectra we observed. We injected a series of secondary signals into our average TRES line profile for TIC 434398831, varying in velocity separation (-50 to 50 $\mathrm{km\,s}^{-1}$), line width (3 to 35 $\mathrm{km\,s}^{-1}$), and primary-secondary flux ratio (0.001 to 0.2). We then attempted to recover each scenario with a double-lined model of the resulting line profiles. We found that for hypothetical sharp lined companions, $v_{\mathrm{rot}}=5\,\mathrm{km\,s}^{-1}$, we can recover signals with $\Delta m\sim 3$ in the optical. Recovery of stellar companions with more significant rotational broadening becomes progressively harder. Companions with rotational broadening similar to that of TIC 434398831 can be recovered at the $\Delta m\sim 1$ level. Figure 11 shows the magnitude limit of our spectroscopic blend analysis as a function of velocity separation from the host star.
@@
To further derive a false-positive probability for any remaining scenarios, we performed a joint analysis combining physically possible binary scenarios from MOLUSC (Wood et al., 2021) with TRICERATOPS (Giacalone et al., 2021). We simulated 100,000 astrophysical false-positive scenarios with MOLUSC for both TIC 434398831 b and c. We made use of radial velocity and $\Delta m$ limits established by the reconnaissance spectroscopic observations and Gaia RUWE and photometry to inform our MOLUSC simulations.
The simulated binaries from MOLUSC and our false-positive analyses were used to rule out TRICERATOPS false-positive scenarios that would not be physically realistic for TIC 434398831. As mentioned above, false-positive scenarios involving a resolved nearby star are ruled out through ground-based transit observations, and our TRES spectroscopy rules out unresolved companions and unresolved background false-positive scenarios within $\Delta\mathord{\sim}3$ magnitudes. The remaining false-positive scenarios identified by TRICERATOPS were diluted transiting planets, resulting from an unresolved background star diluting the light curve of TIC 434398831. This would mean TIC 434398831 b and c were not Neptune-sized planets but Jovian-sized. At the orbital periods of TIC 434398831 b ($P_{b}=3.69$ days) and c ($P_{c}=6.21$ days), the occurrence rates for Jovian-sized planets are $<0.4\%$ (Kunimoto & Matthews, 2020). We therefore confidently present TIC 434398831 b and c as validated planets, as the derived false-positive probabilities for TIC 434398831 b and c are $0.0047_{-0.0013}^{+0.0012}$ and $0.000257_{-0.00022}^{+0.00556}$ respectively.

### v2

@@ 2.3 Ground-based Photometry
We obtained ground-based photometry as part of the TESS Follow-up Observing Program (TFOP; Collins, 2019) [^1] https://tess.mit.edu/followup. We used the TESS Transit Finder, which is a customized version of the Tapir software package (Jensen, 2013), to schedule our transit observations. A summary of our ground-based photometric observations is presented in Table 1, with the light curves shown in Figure 4. The ground-based observations confirm the transits of TIC 434398831 b and c are on target to within 3.1”, and clear the field for nearby eclipsing binary (NEB) scenarios $<2.5$’. We made use of the raw ground-based photometry and simultaneously detrended the transit observations in our global model. Theses observations are made publically available through Exoplanet Follow-up Observing Program (ExoFOP). [^2] https://exofop.ipac.caltech.eduTESStarget.php?id=434398831
@@ 2.3.3 Whitin Observatory
2.4 High-Resolution Imaging
Figure 5: High-resolution image of TIC 434398831 with ’Alopeke on Gemini North. No secondary source was detected. Our $5\sigma$ contrast curves for the red and blue arm are plotted.
To search for visual companions we obtained high-resolution images of TIC 434398831. We observed TIC 434398831 using the red (562 nm) and blue (832 nm) arm of the ‘Alopeke speckle imager. ‘Alopeke is located on the 8-meter Gemini North Telescope (Scott et al., 2021), with a field of view of $2.^{\prime\prime}5\times 2.^{\prime\prime}5$ and a pixel scale of 0.”01 per pixel.
We observed TIC 434398831 with ‘Alopeke on 2024 September 27. Images were reduced using a custom pipeline (Horch et al., 2011; Howell et al., 2011) developed by the ‘Alopeke group. No secondary sources were detected around TIC 434398831. Our observations achieved a contrast ratio of $\Delta$6.63 mag at 0.5” in the red and $\Delta$5.43 mag at 0.5” in the blue. Our resulting $5\sigma$ contrast curves are presented in Figure 5.
@@ 2.5 Reconnasance Spectroscopy
We obtained a set of spectroscopic observations to characterize the spectroscopic properties of the host star and to search for signatures of astrophysical false positive scenarios. We obtained seven epochs of observations with the Tillinghast Reflector Echelle Spectrograph (TRES; Fűrész 2008) on the 1.5 m reflector at the Fred Lawrence Whipple Observatory, Arizona. TRES is a high-resolution echelle spectrograph with a resolving power of $R\approx 44,000$, with a wavelength coverage of $3850-9100\,$Å. Observing strategy and spectral extraction techniques are outlined in Buchhave et al. (2012). For each spectral epoch, a sequence of three consecutive exposures are obtained and combined with sigma clipping applied to mitigate the effects of cosmic ray artefacts. Wavelength calibration is provided via bracketing thorium-argon lamp exposures before and after the target exposure. Spectroscopic atmospheric parameters are derived from each spectrum via the Stellar Parameter Classification (SPC) pipeline as per Buchhave et al. (2012), via a cross correlation between the observed spectrum and a grid of spectral templates from the Kurucz (1992) model atmospheres. An interpolation over the cross correlation function peak heights, as a function of the stellar atmosphere parameters, provides the adopted spectral parameters. Uncertainties in the atmosphere parameters from SPC are estimated based on the scatter in the parameters for a test population that was also classified via the Spectroscopy Made Easy (Valenti & Piskunov, 1996). Relative velocities were derived from a cross-correlation between each epoch spectrum and a template constructed from all TRES observations of the target as per Quinn et al. (2012). Cross-correlation functions are derived over $\sim 25$ of the best orders, least affected by telluric and major Balmer lines. The cross-correlation functions from these orders are added and fit via a Gaussian to derive the relative RV from each epoch. The radial velocity uncertainties are determined from via $\sigma_{\mathrm{rv}}=\mathrm{RMS}/\sqrt{N}$, where RMS is the root mean square scatter of the relative velocities from each order, and $N$ is the number of orders used in the analysis. These are presented in Table 2.
@@ 3.1 TIC 434398831 and Theia 116
Figure 6: Color-Magnitude Diagram (CMD) for candidate Theia 116 members and TIC 434398831 (red star). We plot candidate members with measured rotation periods in orange, filled circles. Candidates without measured rotations are plotted as empty circles.
@@ 3.1.1 Kinematics
TIC 434398831 was identified as a candidate member of Theia 116 in Kounkel & Covey (2019). In brief, Kounkel & Covey (2019) performed a search for comoving stellar populations through analysis of spatial motions from Gaia DR2 (Gaia Collaboration et al., 2018b) via the unsupervised learning algorithm HDBSCAN (McInnes et al., 2017). The ages of identified stellar groups were estimated via fitting an isochrone to the groups’ color-magnitude diagram (see Figure 6). Kounkel & Covey (2019) identified 1108 candidate members of Theia 116. We used Gaia DR3 (Gaia Collaboration et al., 2021) to determine Theia 116 candidate members’ spatial positions and velocities. Figure 7 shows TIC 434398831 relative to the space motion distribution of the Theia 116 group.
@@ 3.1.2 Stellar Rotation
Figure 8: Top panel: Lomb-Scargle periodogram for TIC 434398831 both the average of the full TESS light curve (red) and each sector of TESS observations (grey). We measure a stellar rotation period of $P_{\mathrm{rot}}=1.75\pm 0.12$ days.
@@
Figure 9: Upper panel: Rotation periods for rotationally validated Theia 116 members (orange circles) and TIC 434398831 (red star) as a function of stellar effective temperature. We plot measured rotations of well-characterized clusters for comparison– 40 Myr Tuc-Hor (left, Popinchalk et al., 2023), 80 Myr $\alpha$ Per (middle, Boyle & Bouma, 2023), and 120 Myr Pleiades (right, Rebull et al., 2016a). Lower panel: Lithium 6708Å  doublet equivalent widths (EW) of a subsample of Theia 116 members (orange circles) and TIC 434398831 (red circle). Rotationally validated members of Theia 116 are the filled circles, while empty circles do not exhibit photometric signatures of youth, and are likely contaminating field stars. Literature Li EW’s of Tuc-Hor (Popinchalk et al., 2023), $\alpha$ Per (Balachandran et al., 2011), and Pleiades (Bouvier et al., 2018) members are plotted for comparison. The spread in both rotation and Li abundances of Theia 116 members are consistent with youth. The rotation spread is more consistent with spreads of the pre-zero age main sequence (ZAMS) clusters Tuc-Hor and $\alpha$ Per, than the spread of the ZAMS Pleiades cluster.
@@
We visually examined the full light curve of each candidate member for (1) rotational variability, (2) binarity, (3) non-rotation variability, or (4) no variability. We also checked for aliasing of the true rotation period by phase-folding the light curve at the measured Lomb-Scargle period, half, and double the period. Candidate members are considered rotationally validated if they were classified as rotationally variable stars and we were able to confidently measure a stellar rotation period from the CDIPS light curve. We were able to rotationally validate 224 members of Theia 116, including TIC 434398831, for which we measured a rotation period of $P_{\mathrm{rot}}=1.75\pm 0.12$ days (see Figure 8). We note the beat-like pattern in the TIC 434398831 light curve, corresponding to the additional two peaks surrounding the maximum peak within the periodogram (see Figure 8). These peaks are within 20% of the derived rotation period. While a secondary source could be responsible for the observed beat-like pattern (e.g., Stauffer et al., 2018), the observed features in the light curve and corresponding peaks in the periodogram are fully consistent with differential rotation of the young star (Aigrain et al., 2015; Rebull et al., 2016b; Santos et al., 2017; Stauffer et al., 2018). Our measured rotation periods for TIC 434398831 and Theia 116 members are shown in Figure 9.
@@ 3.1.3 Lithium Depletion
We measured the equivalent width of the lithium doublet at 6708 Å as per Zhou et al. (2021). We first performed a least-squares deconvolution analysis to derive the line profile of each star, and thereby measure its rotational velocity. We then fit the spectral region around the 6708 Å line via a set of Gaussian profiles. We assumed the two lines of the lithium doublet have the same amplitude and the same width as that of the stellar rotational broadening velocity. We also simultaneously modeled the nearby Fe I line at 6707.43 Å to account for its contamination to the lithium doublet line strength, adopting the same broadening width, but allowing its amplitude to be free. We adopted the integral below our model of the lithium doublet as the equivalent width of the lithium line. All spectra were checked for signatures of binarity, no members exhibited double-line profiles. Effective temperatures were also derived from each spectrum via SPC as per Section 2.5. The measured lithium abundances are presented in Table 4. We compare our measured lithium equivalent widths against literature values for well-characterized clusters in Figure 9.
@@ 3.1.4 Isochrone modeling of membership SEDs
We fitted for the cluster age, extinction, and stellar mass of each rotationally validated member.These parameters were then interpolated onto a MESA Isochrones and Stellar Tracks (MIST) isochrone (Dotter, 2016) interpolated with the minimint code (Koposov, 2021). We adopt the absolute magnitudes that are provided by the MIST models and compute apparent magnitudes with the incorporation of the stars’ known parallax and modeled extinction. We assumed solar metallicity and Gaia parallax values. We compared estimated magnitudes to observed magnitudes from the Gaia DR3 G, BP, and RP bands, WISE W1, W2, and W3 bands, and 2MASS J, H, and Ks bands. The SED of TIC 434398831 is presented in Figure 10. The best-fit cluster age and its associated uncertainties were explored via an MCMC implemented in emcee. We estimated a SED cluster age of 55$\pm$4 Myr.
@@ 3.1.5 Age of Theia 116
We attempted numerous joint age analyses that incorporated lithium distributions, isochrones, and rotation distributions together for Theia 116, but derived age posteriors with uncertainties too small to be realistic ($<3$ Myr), likely resulting from intrinsic uncertainties in stellar models unaccounted for in our analysis. For the remainder of this work, we therefore adopt a more conservative cluster age of $\pm 6$ Myr, which encompasses the posteriors of the age analyses from independent lithium distribution modeling and the isochrone modeling. We present our age posteriors from the stellar rotation, lithium abundance, and SED analyses in Figure 11.
@@ 3.2 Stellar flare rate of TIC 434398831
In addition to heightened photometric variability induced by rapid rotation, stellar flares from active young stars bring additional challenges when searching for transit planets (e.g. AU Mic c Plavchan et al., 2020). At $\mathord{\sim}61$ Myr, young G-type stars are still expected to have a heightened stellar flare rate when compared to their gigayear old counterparts (Feinstein et al., 2024). Visual examination of the TESS light curves showed evidence for stellar flares detected in the 10 min FFI observations. However, most flare identification algorithms are tuned for TESS 20 second or 2 minute cadence data (e.g. Feinstein et al., 2020b).
Following Howard (2022), we identified flares as deviations $\geq 4.5\sigma$ above the detrended TESS light curve. From TESS 30-minute and 10-minute observations in Sectors 6, 33, 43, 44, and 45, we identified 9 possible flares. These possible events were visually vetted, excluding detected deviations that occur at the edge of the TESS orbit. This resulted in 8 flare events across the 5 sectors of TESS observations. We derived a stellar flare rate of $2.0\times 10^{-2}$ flares per day for TIC 434398831, slightly lower than expected stellar flare rates ($\mathord{\sim}10^{-1}$ flares per day) for $>50$ Myr G-type stars (Feinstein et al., 2024). It is possible the actual flare rate is higher, as stellar flares can occur on time scales shorter than 30 and/or 10 minutes. We repeated this exercise using the 200 s data from sectors 71 and 72 which are likely to be more sensitive to flare events due to the shorter cadence (Howard & MacGregor, 2022). We identified 4 flare events (5 excursions from the detrended TESS light curve, 1 excluded event due to its location near the edge of the TESS orbit), and measured a flare rate of $\times 10^{-1}$ flares per day in the 200 s data.
We repeated our flare rate calculation using the flare identification criteria implemented in stella (Feinstein et al., 2020a), a convoluted neural network trained to detect flare events in the short cadence TESS data, to identify possible flare events in the 200 s TESS observations from Sectors 71 and 72. stella requires candidate flare events to have an amplitude $\geq 1.5\sigma$ above the detrended light curve, the two cadences immediately following the candidate event to have a $\geq 1\sigma$ deviation above the detrended light curve as well, the cadences immediately before and after the candidate event must have lower amplitudes, and the amplitude of the candidate flare must be $\geq 0.5\%$. Three candidate flare events within the 200s data meet this criterion. We note that due to the long cadence data, we were unable to use stella directly on the TESS data for TIC 434398831. This gives an estimated flare rate of 7.6×$\times 10^{-2}$ flares per day, which aligns with measured flare rates for young G dwarf stars with stella as per Feinstein et al. (2024). The true flare rate may still be higher, as the apparent flare maximum can be reduced in the longer TESS cadences.
@@ 5 Investigating possible false-positive scenarios
Due to the large size of TESS’s pixels, some transit signals that pass vetting procedures may result from false positives. As TIC 434398831 b and c were not released as a TOI, we investigate possible instrumental and astrophysical false-positive scenarios to validate the planets. The transit detections of both TIC 434398831 b and c by multiple instruments ruled out instrumental artifact false-positive scenarios. Spectroscopic follow-up observations ruled out TIC 434398831being an eclipsing binary.
@@
where $t_{12}$ is the ingress duration, $t_{13}$ is the time between first and third contact, and $\delta$ is the transit depth. An unresolved star would have to be within $\Delta m_{\mathrm{\textit{TESS}}}\lesssim 3.16$ and $\Delta m_{\mathrm{\textit{TESS}}}\lesssim 2.48$ (at a $3\sigma$ upper limit) to produce the signals corresponding to the transits of TIC 434398831 b and c respectively. We estimated the density of stars within $\Delta m_{\mathrm{\textit{TESS}}}\leq 3.16$ in the general area of TIC 434398831 to determine the likelihood of an unresolved background star being the source of the transit signals. We queried Gaia for resolved stars within a 1 arcmin cone centered around TIC 434398831, which yielded 32 stars. We cross-matched the Gaia sources with the TIC to calculate $\Delta m_{TESS}$. We derive a density of $4.42\times 10^{-5}$ stars per square arcsecond within $\Delta m_{\mathrm{\textit{TESS}}}\leq 3.16$ of TIC 434398831.
We also made use of our TRES spectroscopic observations to help establish the magnitude limits of hypothetical stellar neighbors that may be blended with our target star. Following Zhou et al. (2021), we used the line profiles derived from our TRES observations to determine if scenarios involving non-associated stellar companions with differing radial velocities can be ruled out, as they would yield a detectable set of lines in the spectra we observed. We injected a series of secondary signals into our average TRES line profile for TIC 434398831, varying in velocity separation (-50 to 50 $\mathrm{km\,s}^{-1}$), line width (3 to 35 $\mathrm{km\,s}^{-1}$), and primary-secondary flux ratio (0.001 to 0.2). We then attempted to recover each scenario with a double-lined model of the resulting line profiles. We found that for hypothetical sharp lined companions, $v_{\mathrm{rot}}=5\,\mathrm{km\,s}^{-1}$, we can recover signals with $\Delta m\sim 3$ in the optical if the two stellar components are separated in velocity by $>10\mathrm{km\,s}^{-1}$. Recovery of stellar companions with more significant rotational broadening becomes progressively harder. Companions with rotational broadening similar to that of TIC 434398831 can be recovered at the $\Delta m\sim 1$ level. Figure 12 shows the magnitude limit of our spectroscopic blend analysis as a function of velocity separation from the host star.
@@
To further derive a false-positive probability for any remaining scenarios, we performed a joint analysis combining physically possible binary scenarios from MOLUSC (Wood et al., 2021) with TRICERATOPS (Giacalone et al., 2021). We simulated 100,000 possible binary configurations between TIC434398831 and an unseen companion(s) with MOLUSC for both TIC 434398831 b and c. We made use of radial velocity, HR-contrast curves, Gaia RUWE and photometry to inform our MOLUSC simulations.
The simulated binaries from MOLUSC and our false-positive analyses were used to rule out TRICERATOPS false-positive scenarios (e.g., nearby eclipsing binaries, nearby transiting planets, dilution of the light curve due to an unresolved background star) that would not be physically realistic for TIC 434398831. As mentioned above, false-positive scenarios involving a resolved nearby star are ruled out through ground-based transit and HR-imaging observations. The remaining false-positive scenarios identified by TRICERATOPS were diluted transiting planets, resulting from an unresolved background star diluting the light curve of TIC 434398831. This would mean TIC 434398831 b and c were not Neptune-sized planets but Jovian-sized. At the orbital periods of TIC 434398831 b ($P_{b}=3.69$ days) and c ($P_{c}=6.21$ days), the occurrence rates for Jovian-sized planets are $<0.4\%$ (Kunimoto & Matthews, 2020). We therefore confidently present TIC 434398831 b and c as validated planets, as the derived false-positive probabilities for TIC 434398831 b and c are $0.0047_{-0.0013}^{+0.0012}$ and $0.000257_{-0.00022}^{+0.00556}$ respectively.